ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

11. Mathematical Physics

1991 MS Classification: 17, 81 Bytsko, Andrei, Steklov Mathematics Institute, St. Petersburg, Russia On factorized Virasoro characters. The characters of minimal Virasoro models are known explicitly in an infinite sum form. I will present some cases in which characters of minimal models or their linear combinations are factorizable, i. e. admit an infinite product form, or more precisely, are given as a product of finite number of functions known in mathematical physics as the quantum dilogarithm. The factorized form of the characters significantly simplifies the task of finding various types of identities relating characters from different sectors and different models. The characters of some super-conformal models will also appear in this context. References: A. G. Bytsko and A. Fring, Anyonic interpretation of Virasoro characters and the thermodynamic Bethe ansatz, to appear in Nucl. Phys. B 521 (1998); A. G. Bytsko and A. Fring, On factorized Virasoro characters in minimal models, in preparation.